# A Survey on LLM-as-a-Judge

### v1

@@ Abstract
Accurate and consistent evaluation is crucial for decision-making across numerous fields, yet it remains a challenging task due to inherent subjectivity, variability, and scale. Large Language Models (LLMs) have achieved remarkable success across diverse domains, leading to the emergence of ”LLM-as-a-Judge,” where LLMs are employed as evaluators for complex tasks. With their ability to process diverse data types and provide scalable, cost-effective, and consistent assessments, LLMs present a compelling alternative to traditional expert-driven evaluations. However, ensuring the reliability of LLM-as-a-Judge systems remains a significant challenge that requires careful design and standardization. This paper provides a comprehensive survey of LLM-as-a-Judge, addressing the core question: How can reliable LLM-as-a-Judge systems be built? We explore strategies to enhance reliability, including improving consistency, mitigating biases, and adapting to diverse assessment scenarios. Additionally, we propose methodologies for evaluating the reliability of LLM-as-a-Judge systems, supported by a novel benchmark designed for this purpose. To advance the development and real-world deployment of LLM-as-a-Judge systems, we also discussed practical applications, challenges, and future directions. This survey serves as a foundational reference for researchers and practitioners in this rapidly evolving field. https://github.com/IDEA-FinAI/LLM-as-Evaluator.
@@ 1. Introduction
Recently, Large Language Models (LLMs) have achieved remarkable success in numerous domains, ranging from artificial intelligence and software engineering to education and social science. The adoption of LLMs as evaluators—commonly referred to as ”LLM-as-a-Judge” (Zheng et al., 2023b)—has surged, driven by their ability to emulate human-like reasoning and decision-making processes. This capability enables LLMs to undertake roles traditionally reserved for human experts, offering a cost-effective and scalable alternative. For instance, we usually rely on experts to evaluate the accuracy of mathematics and physics competition questions at the Olympiad level (He et al., 2024a), which can be assessed through LLM-as-a-Judge now. Additionally, in recent peer reviews of research submissions, LLM-as-a-Judge is introduced to address the rising number of paper submissions and reviewer workload [^1] https://blog.iclr.cc/2024/10/09/iclr2025-assisting-reviewers/, which is designed to identify potential issues in reviews and offer constructive feedback to reviewers. These trends underscore a key motivation for adopting LLM-as-a-Judge: the potential to enhance evaluation efficiency while addressing limitations inherent in human assessments, such as scalability and consistency.
LLM-as-a-Judge presents a compelling alternative to both human evaluations and traditional automated methods, offering distinct advantages in scalability, efficiency, and adaptability. Human evaluations, though often considered the gold standard, face challenges in scalability, cost, and consistency. They are time-consuming, require substantial expert effort, and are expensive to scale due to limited availability of qualified evaluators. Coordinating and training evaluators adds complexity, and fatigue during lengthy tasks can compromise reliability and accuracy. In contrast, LLMs provide scalable, cost-effective, and efficient evaluations with reduced subjective variability, enhancing objectivity. Traditional automated methods, such as BLEU, ROUGE, and METEOR for software artifact summarization, often fail to align with human judgment or provide clear insights in specialized domains like software engineering. LLMs offer flexibility to process diverse input types, including text, semi-structured data, and multi-modal content, allowing evaluations to integrate qualitative insights with quantitative rigor. This human-aligned adaptability makes LLMs effective for complex, context-aware assessments beyond the limits of conventional metrics (Wang et al., 2023d; Li et al., 2023b; Chen et al., 2024c).
Despite its wide advantages, LLM-as-a-Judge poses significant challenges for reliability. This necessitates the capability of LLM-as-a-Judge framework to subsume under rules, that is, to distinguish whether something falls under a given rule (Kant, 1781). As LLM-as-a-Judge becomes more commonly used as an effective evaluator in different areas, collecting evaluations with LLM-as-a-Judge is relatively simple. Therefore, central to this survey is the fundamental question: How to build reliable LLM-as-a-Judge systems? To address this question, we explore two core aspects: (1) strategies for enhancing the reliability of LLM-as-a-Judge and (2) methodologies for evaluating reliability of LLM-as-a-Judge systems themselves.
@@
For the first aspect of enhancing LLM-as-a-Judge reliability, we review the main strategies aimed at optimizing their performance for diverse evaluation tasks. These strategies include improving consistency, mitigating biases, and refining adaptability to different assessment scenarios. For the second aspect, we examine the metrics, datasets, and methodologies used to evaluate the performance of LLM-as-a-Judge systems, discussing potential sources of bias and corresponding mitigation techniques. Building on this foundation, we introduce a novel benchmark specifically designed for LLM-as-a-Judge evaluations. Using established metrics and datasets, this benchmark provides a framework for analyzing the effectiveness of various reliability enhancement strategies. Additionally, we explore practical application scenarios, identify specific challenges unique to each context, and propose solutions to address these issues. Finally, we discuss future research directions, emphasizing key areas for advancing the reliability, scalability, and applicability of LLM-as-a-Judge systems.
This paper aims to provide a comprehensive overview of the LLM-as-a-Judge research landscape while offering insights into how reliable LLM-as-a-Judge can be constructed. We hope that this work will serve as a valuable reference for researchers and practitioners, fostering further research and facilitating the real-world deployment of LLM-as-a-Judge. The rest of this survey is organized as Figure 1. In Section 2, we provide a comprehensive overview of the LLM-as-a-Judge field. We define LLM-as-a-Judge through formal and informal definitions, and categorize existing methods and approaches for its use. For a quick guide on implementing an LLM-as-a-Judge for specific scenarios, you can find answers in Quick Practice (2.5). Then, we discuss the problems of ”How to improve” and ”how to evaluate” LLM-as-a-Judge in n Section 3 to 4 . Next, we show the applications of LLM-as-a-Judge in Section 6. The discussions of the challenges and future directions come at last in Section 7 and Section 8.
Figure 1. Paper Structure.
@@ 2. Background and Method
In evaluative tasks, especially those that are subjective, human assessment is often considered the gold standard due to its reliable and open-ended nature (Shi et al., 2024; Gao et al., 2023). However, this approach is typically slow and costly (Zheng et al., 2023b; Zhu et al., 2023a; Yuan et al., 2023). To address these challenges, LLMs are increasingly employed as substitutes for human evaluators. Since these models are frequently trained using Reinforcement Learning from Human Feedback (RLHF), they demonstrate strong alignment with human perspectives, leading to the approach known as ”LLM-as-a-Judge”.
In general, LLM-as-a-Judge is a process that uses LLM to evaluate different objects in various scenarios for diverse tasks. For instance, roles such as ”Assessors”, ”Critics,” and ”Verifiers” utilize LLMs to facilitate evaluation at different stages of the process, whether during intermediate steps or throughout the entire workflow. To date, definition of how to effectively use LLM-as-a-Judge for evaluation tasks has been largely informal or vague, lacking clear and formal expression. Therefore, we will provide a formal definition of LLM-as-Evaluator as follows:
@@
The formulation of LLM-as-a-Judge reflects that LLM is a type of auto-regressive generative model, which generates subsequent content based on the context then obtain target evaluation from it. The form of LLM-as-a-Judge illustrates how we utilize LLM for evaluation tasks, encompassing input design, model selection and training, as well as output post-processing. The different basic approaches of implementing LLM-as-a-Judge can be classified according to the formulation: In-Context Learning, Model Selection, Post-processing Method and Evaluation Pipeline, which concluded in Figure 2. By following this pipeline, we can build a basic LLM-as-a-Judge for evaluation. A faster practice guide is available in section 2.5.
@@ 2.1. In-Context Learning
To apply LLM-as-a-Judge, it is helpful to start by defining the evaluation task using In-Context Learning methods. This process involves two key aspects: the design of prompt and input. For input design, it is important to consider the type of variables to be evaluated (such as text, image, or video), the manner of input (e.g., individually, in pairs, or in batches), and the position of the input (e.g., at the beginning, middle, or end). As for the prompt design, four different methods can be adopted, as illustrated in Figure 2. The four methods include generating scores, solving true/false questions, conducting pairwise comparisons, and making multiple-choice selections. Further details will be provided in the following sections.
@@ 2.1.1. Generating scores
It is quite intuitive to represent an evaluation using a corresponding score. What requires more careful consideration, however, is the nature and range of the score used for evaluation. The score can be discrete, with common ranges like 1-3, 1-5 (Jones et al., 2024), or 1-10 (Zhu et al., 2023a; Li et al., 2023b). Alternatively, it can be continuous, ranging from 0 to 1 or 0 to 100 (Xiong et al., 2024). The simplest way to score is through the context, setting the range of scores and the main criteria for scoring. For example, ”Please rate the helpfulness, relevance, accuracy, level of details of their responses. Each assistant receives an overall score on a scale of 1 to 10, where a higher score indicates better overall performance” (Zhu et al., 2023a). A slightly more complex way is to provide more detailed scoring criteria. More complex scoring situations can be as Language-Model-as-an-Examiner (Bai et al., 2024a), which use Likert scale scoring functions as an absolute evaluative measure showed in Figure 3. The evaluator assigns scores to a given response along predefined dimensions including accuracy, coherence, factuality and comprehensiveness. Each of these dimensions is scored on a scale of 1 to 3, ranging from worst to best. The evaluator is also asked to provide an overall score ranging from 1 to 5, based on the scores assigned to the previous 4 dimensions. This score serves as an indicator of the overall quality of the answer.
@@ 2.1.2. Solving Yes/No questions
This type of evaluation is often utilized in intermediate processes, creating the conditions for a feedback loop. For example, it promotes a self-optimization cycle, as seen in Reflexion (Shinn et al., 2024), which generates verbal self-reflections to provide valuable feedback for future attempts. In scenarios with sparse reward signals, such as a binary success status (success/fail), the self-reflection model uses the current trajectory and persistent memory to generate nuanced and specific feedback. Similarly, in self-improvement contexts (Tian et al., 2023b), Yes/No questions can be employed to evaluate custom phrases, such as "Modification needed." and "No modification needed.", facilitating entry into the next cycle. Moreover, these evaluations are common for testing knowledge accuracy and assessing whether statements align with established facts (Sun et al., 2023c), like ”Given a question and the associated retrieved knowledge graph triples (entity, relation, entity), you are asked to answer whether it’s sufficient for you to answer the question with these triples and your knowledge (Yes or No).” A detailed and specific example can be seen in the Figure 4.
@@ 2.1.3. Conducting pairwise comparisons
Pairwise comparison is a well-established method that has significantly impacted a variety of fields (Qin et al., 2023). As noted by (Liu et al., 2024c), LLM and human evaluations are more aligned in the context of pairwise comparisons compared to score-based assessments. Numerous studies have demonstrated that pairwise comparative assessments outperform other judging methods in terms of positional consistency (Zheng et al., 2023c; Liusie et al., 2024). Furthermore, pairwise comparisons can be extended to more complex relation-based assessment frameworks, such as list-wise comparisons, using advanced ranking algorithms (Qin et al., 2023; Liu et al., 2024c), data filtering (Yuan et al., 2023). In pairwise comparative assessments, LLM-as-a-Judge is prompted to select the response that better answers the question at hand. To accommodate the possibility of a tie, several option modes are introduced. The Two-Option mode requires judges to choose the better response from two given options. The Three-Option mode introduces an additional choice, allowing judges to indicate a tie if neither response is preferable. Evaluations typically involve determining the outcomes of win, tie, or loss for responses (Wang et al., 2023d) through pairwise comparisons, with win rounds counted for each response. The Four-Option mode further expands the choices, allowing judges to classify responses as either a ”both good tie” or a ”both bad tie.”
@@ 2.2.2. Fine-tuned LLM
However, relying on external API for evaluation may introduce consideration about privacy leakage, and the opacity of API models also challenges the evaluation reproducibility. Therefore, follow-up works suggest fine-tuning language models specialized in evaluations. For instance, PandaLM (Wang et al., 2023d) constructs data based on Alpaca instructions and GPT-3.5 annotation, and then fine-tunes LLaMA-7B (Touvron et al., 2023a) as an evaluator model. JudgeLM (Zhu et al., 2023a) constructs data from diversified instruction sets and GPT-4 annotations, and fine-tunes Vicuna (Touvron et al., 2023b) as a scalable evaluator model. Auto-J (Li et al., 2023b) constructs evaluation data upon multiple scenarios to train a generative evaluator model, which can provide both evaluation and critical opinion. Prometheus (Kim et al., 2023) defines thousands of evaluation criteria and construct a feedback dataset based on GPT-4, and fine-tunes a fine-grained evaluator model. The typical process for fine-tuning a judge model involves three main steps. Step 1: Data Collection. The training data generally consists of three components: instructions, the objects to be evaluated, and evaluations. Instructions are typically sourced from instruction datasets, while evaluations can come from either GPT-4 or human annotations. Step 2-Prompt Design. The structure of the prompt template can vary based on the evaluation scheme, which already detailed in § 2.1. Step 3: Model Fine-Tuning. Using the designed prompts and collected data, the fine-tuning process for the evaluator model typically adheres to the instruction fine-tuning paradigm (Ouyang et al., 2022b). The model receives an instruction along with one or more responses to generate output that includes evaluation results and possibly explanations.
After fine-tuning, the evaluator model can be employed to evaluate the target object. While these fine-tuned models often demonstrate superior performance on self-designed test sets, they are identified several limitations in their evaluation capabilities, which detailed in Section 4.2.

### v2

@@ Abstract
Accurate and consistent evaluation is crucial for decision-making across numerous fields, yet it remains a challenging task due to inherent subjectivity, variability, and scale. Large Language Models (LLMs) have achieved remarkable success across diverse domains, leading to the emergence of ”LLM-as-a-Judge,” where LLMs are employed as evaluators for complex tasks. With their ability to process diverse data types and provide scalable, cost-effective, and consistent assessments, LLMs present a compelling alternative to traditional expert-driven evaluations. However, ensuring the reliability of LLM-as-a-Judge systems remains a significant challenge that requires careful design and standardization. This paper provides a comprehensive survey of LLM-as-a-Judge, addressing the core question: How can reliable LLM-as-a-Judge systems be built? We explore strategies to enhance reliability, including improving consistency, mitigating biases, and adapting to diverse assessment scenarios. Additionally, we propose methodologies for evaluating the reliability of LLM-as-a-Judge systems, supported by a novel benchmark designed for this purpose. To advance the development and real-world deployment of LLM-as-a-Judge systems, we also discussed practical applications, challenges, and future directions. This survey serves as a foundational reference for researchers and practitioners in this rapidly evolving field. The associated resources can be accessed at https://github.com/IDEA-FinAI/LLM-as-a-Judge.
@@ 1. Introduction
Recently, Large Language Models (LLMs) have achieved remarkable success across numerous domains, ranging from technical fields to the humanities and social sciences. Building on their success, the concept of using LLMs as evaluators—commonly referred to as ”LLM-as-a-Judge” (Zheng et al., 2023b)—has gained significant attention, where LLMs are tasked with determining whether something falls within the scope of a given rule. This growing interest stems from LLMs’ ability to mimic human-like reasoning and thinking processes, enabling them to take on roles traditionally reserved for human experts while offering a cost-effective solution that can be effortlessly scaled to meet increasing evaluation demands. For instance, employing LLM-as-a-Judge in the academic peer-review [^1] https://blog.iclr.cc/2024/10/09/iclr2025-assisting-reviewers/ process can help handle the rapid increase in submissions while maintaining expert-level judgment.
Before the era of LLMs, achieving a balance between comprehensive and scalable evaluation had posed a persistent challenge. On the one hand, commonly used subjective methods like expert-driven assessments integrate holistic reasoning and nuanced contextual understanding, making them the gold standard in comprehensiveness  (Shi et al., 2024a; Gao et al., 2023). However, these approaches are expensive, challenging to scale, and susceptible to inconsistency. On the other hand, objective assessment methods like automatic metrics offer strong scalability and consistency. For example, tools such as BLEU (Papineni et al., 2002) or ROUGE (Lin, 2004) can quickly score machine-generated translations or summaries against reference texts without human intervention. Yet these metrics, which heavily rely on surface-level lexical overlaps, often fail to capture deeper nuances of meaning, coherence, or logical reasoning. As a result, they perform poorly when evaluating tasks like open-ended story generation or complex instructional texts (Schluter, 2017). With the rise of advanced LLM techniques, the “LLM-as-a-Judge” has emerged as a potential idea to combine the strengths of above two ways of evaluation. LLMs have been proven to merge the scalability of automatic methods with the nuanced, context-sensitive reasoning found in expert judgments (Zheng et al., 2023b; Wang et al., 2023d; Zhu et al., 2023a; Li et al., 2023b; Chen et al., 2024d). Moreover, LLMs might become flexible enough to handle multimodal inputs citemllm-as-a-judge under decent prompt learning or fine-tuning (Khattak et al., 2023), which is a unique advantage compared with traditional methods. These advantages suggest that the LLM-as-a-Judge approach could serve as a novel and broadly applicable paradigm for addressing complex and open-ended evaluation problems.
While LLM-as-a-Judge holds significant potential as a scalable and adaptable evaluation framework compared to expert-driven human evaluations and traditional automated metrics (Wang et al., 2024a), its adoption is hindered by two key challenges. The first challenge is the absence of a systematic review, which reflects the lack of formal definitions, fragmented understanding, and casual usage practices in the study of LLM-as-a-Judge. These gaps make it difficult for researchers and practitioners to fully understand LLM-as-a-Judge and apply it effectively. Building on this, the second challenge involves addressing ongoing concerns about reliability (Yu et al., 2024b), as merely employing LLM-as-a-Judge does not guarantee evaluations that are both accurate and aligned with established standards. These challenges demand not only a deeper assessment of the outputs generated by LLM-as-a-Judge but also a crucial investigation into the question: How to build reliable LLM-as-a-Judge systems?
@@
To address these challenges, this paper provides a systematic review of research on LLM-as-a-Judge, offering a comprehensive overview of the field while exploring strategies for building reliable LLM-as-a-Judge systems. We begin by defining LLM-as-a-Judge through both formal and informal definitions, answering the foundational query: ’What is LLM-as-a-Judge?’ Next, we categorize existing methods and approaches for its use, exploring how LLM-as-a-Judge can be effectively implemented in practice. Following this, we tackle the critical question: How to build reliable LLM-as-a-Judge systems? To answer this, we explore two core aspects: (1) strategies to enhance the reliability of LLM-as-a-Judge systems and (2) methodologies for evaluating the reliability of these systems. For the first aspect, we review key strategies to optimize the performance of LLM-as-a-Judge for diverse evaluation tasks. For the second aspect, we examine the metrics, datasets, and methodologies used to evaluate the performance of LLM-as-a-Judge systems, highlighting potential sources of bias and methods for their mitigation. Building on this foundation, we propose a novel benchmark specifically designed for evaluating LLM-as-a-Judge systems. Additionally, we explore practical application scenarios, identify challenges unique to each context, and propose solutions to address these issues. Finally, we discuss future research directions, emphasizing key areas for improving the reliability, scalability, and applicability of LLM-as-a-Judge systems.
The rest of this survey is organized as Figure 1. Section 2 provides an overview of the LLM-as-a-Judge field, including its definitions and categorization of existing methods.For a quick guide on implementing an LLM-as-a-Judge for specific scenarios, you can find answers in Quick Practice (2.5). Strategies for enhancing and evaluating the reliability of LLM-as-a-Judge systems are discussed in Sections 3 and 4, respectively. Section 6 explores practical applications, while Sections 7 and 8 address challenges and outline future research directions. Finally, Section 9 presents our conclusions.
Figure 1. The overall framework of this paper.
@@ 2. Background and Method
The ability of LLMs to mimic human reasoning and evaluate specific inputs against a set of predefined rules has paved the way for ”LLM-as-a-Judge.” Their scalability, adaptability, and cost-effectiveness make them well-suited for managing a growing number of evaluative tasks that were traditionally done by humans. These abilities are key in utilizing LLMs flexibly across various evaluation scenarios and objectives. The adoption of LLM-as-a-Judge has progressed rapidly. Initially, the primary focus of LLMs was on language generation and comprehension. With advancements in training paradigms like Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022a), LLMs became more aligned with human values and reasoning processes. This alignment has allowed LLMs to transition from generative tasks to evaluative roles. At its core, LLM-as-a-Judge refers to the use of LLMs to evaluate objects, actions, or decisions based on predefined rules, criteria, or preferences. It encompasses a broad spectrum of roles, including: Graders (Trung et al., 2024), Evaluators/Assessors (Li et al., 2024b; Zhang et al., 2024b), Critics (Ke et al., 2024; Xiong et al., 2024; Putta et al., 2024), Verifiers (Ling et al., 2024; Shinn et al., 2023; Wang et al., 2024c), Examiners (Bai et al., 2023a), Reward/Ranking Models (Shi et al., 2024b; Sun et al., 2023a; Luo et al., 2023a; Yuan et al., 2023), etc.
Currently, the definition of how to effectively use LLM-as-a-Judge for evaluation tasks is largely informal or vague, lacking clear and formal expression. Therefore, we will start with a formal definition of LLM-as-Evaluator as follows:
@@
The formulation of LLM-as-a-Judge reflects that LLM is a type of auto-regressive generative model, which generates subsequent content based on the context then obtain target evaluation from it. It illustrates how we utilize LLM for evaluation tasks, encompassing input design, model selection and training, as well as output post-processing. The basic approaches of implementing LLM-as-a-Judge can be classified according to the formulation: In-Context Learning, Model Selection, Post-processing Method and Evaluation Pipeline, which concluded in Figure 2. By following this pipeline, one can build a basic LLM-as-a-Judge for evaluation. A quick practice guide is available in section 2.5.
@@ 2.1. In-Context Learning
To apply LLM-as-a-Judge, evaluation tasks are typically defined using In-Context Learning methods, which provide instructions and examples to guide the model’s reasoning and judgment. This process involves two key aspects: the design of prompt and input. For input design, it is important to consider the type of variables to be evaluated (such as text, image, or video), the manner of input (e.g., individually, in pairs, or in batches), and the position of the input (e.g., at the beginning, middle, or end). As for the prompt design, four different methods can be adopted, as illustrated in Figure 2. The four methods include generating scores, solving true/false questions, conducting pairwise comparisons, and making multiple-choice selections. Further details will be provided in the following sections.
@@ 2.1.1. Generating scores
It is quite intuitive to represent an evaluation using a corresponding score. What requires more careful consideration, however, is the nature and range of the score used for evaluation. The score can be discrete, with common ranges like 1-3, 1-5 (Jones et al., 2024), or 1-10 (Zhu et al., 2023a; Li et al., 2023b). Alternatively, it can be continuous, ranging from 0 to 1 or 0 to 100 (Xiong et al., 2024). The simplest way to score is through the context, setting the range of scores and the main criteria for scoring. For example, ”Please rate the helpfulness, relevance, accuracy, level of details of their responses. Each assistant receives an overall score on a scale of 1 to 10, where a higher score indicates better overall performance” (Zhu et al., 2023a). A slightly more complex way is to provide more detailed scoring criteria. More complex scoring situations can be as Language-Model-as-an-Examiner (Bai et al., 2023a), which use Likert scale scoring functions as an absolute evaluative measure showed in Figure 3. The evaluator assigns scores to a given response along predefined dimensions including accuracy, coherence, factuality and comprehensiveness. Each of these dimensions is scored on a scale of 1 to 3, ranging from worst to best. The evaluator is also asked to provide an overall score ranging from 1 to 5, based on the scores assigned to the previous 4 dimensions. This score serves as an indicator of the overall quality of the answer.
@@ 2.1.2. Solving Yes/No questions
This type of evaluation is often utilized in intermediate processes, creating the conditions for a feedback loop. For example, it promotes a self-optimization cycle, as seen in Reflexion (Shinn et al., 2023), which generates verbal self-reflections to provide valuable feedback for future attempts. In scenarios with sparse reward signals, such as a binary success status (success/fail), the self-reflection model uses the current trajectory and persistent memory to generate nuanced and specific feedback. Similarly, in self-improvement contexts (Tian et al., 2024), Yes/No questions can be employed to evaluate custom phrases, such as "Modification needed." and "No modification needed.", facilitating entry into the next cycle. Moreover, these evaluations are common for testing knowledge accuracy and assessing whether statements align with established facts (Sun et al., 2023d), like ”Given a question and the associated retrieved knowledge graph triples (entity, relation, entity), you are asked to answer whether it’s sufficient for you to answer the question with these triples and your knowledge (Yes or No).” A detailed and specific example can be seen in the Figure 4.
@@ 2.1.3. Conducting pairwise comparisons
Pairwise comparison is a well-established method that has significantly impacted a variety of fields (Qin et al., 2024). As noted by (Liu et al., 2024b), LLM and human evaluations are more aligned in the context of pairwise comparisons compared to score-based assessments. Numerous studies have demonstrated that pairwise comparative assessments outperform other judging methods in terms of positional consistency (Zheng et al., 2023c; Liusie et al., 2024). Furthermore, pairwise comparisons can be extended to more complex relation-based assessment frameworks, such as list-wise comparisons, using advanced ranking algorithms (Qin et al., 2024; Liu et al., 2024b), data filtering (Yuan et al., 2023). In pairwise comparative assessments, LLM-as-a-Judge is prompted to select the response that better answers the question at hand. To accommodate the possibility of a tie, several option modes are introduced. The Two-Option mode requires judges to choose the better response from two given options. The Three-Option mode introduces an additional choice, allowing judges to indicate a tie if neither response is preferable. Evaluations typically involve determining the outcomes of win, tie, or loss for responses (Wang et al., 2023d) through pairwise comparisons, with win rounds counted for each response. The Four-Option mode further expands the choices, allowing judges to classify responses as either a ”both good tie” or a ”both bad tie.”
@@ 2.2.2. Fine-tuned LLM
However, relying on external API for evaluation may introduce consideration about privacy leakage, and the opacity of API models also challenges the evaluation reproducibility. Therefore, subsequent studies recommend refining language models tailored for evaluations by emphasizing the use of pairwise comparisons or grading. For instance, PandaLM (Wang et al., 2023d) constructs data based on Alpaca instructions and GPT-3.5 annotation, and then fine-tunes LLaMA-7B (Touvron et al., 2023a) as an evaluator model. JudgeLM (Zhu et al., 2023a) constructs data from diversified instruction sets and GPT-4 annotations, and fine-tunes Vicuna (Touvron et al., 2023b) as a scalable evaluator model. Auto-J (Li et al., 2023b) constructs evaluation data upon multiple scenarios to train a generative evaluator model, which can provide both evaluation and critical opinion. Prometheus (Kim et al., 2023) defines thousands of evaluation criteria and construct a feedback dataset based on GPT-4, and fine-tunes a fine-grained evaluator model.
The typical process for fine-tuning a judge model involves three main steps. Step 1: Data Collection. The training data generally consists of three components: instructions, the objects to be evaluated, and evaluations. Instructions are typically sourced from instruction datasets, while evaluations can come from either GPT-4 or human annotations. Step 2-Prompt Design. The structure of the prompt template can vary based on the evaluation scheme, which already detailed in § 2.1. Step 3: Model Fine-Tuning. Using the designed prompts and collected data, the fine-tuning process for the evaluator model typically adheres to the instruction fine-tuning paradigm (Ouyang et al., 2022b). The model receives an instruction along with one or more responses to generate output that includes evaluation results and possibly explanations.
After fine-tuning, the evaluator model can be employed to evaluate the target object. While these fine-tuned models often demonstrate superior performance on self-designed test sets, they are identified several limitations in their evaluation capabilities,which detailed in Section 4.2. The current prompt and fine-tuning dataset designs often result in evaluation LLMs with poor generalization, making them difficult to compare with strong LLMs like GPT-4.